### v1

# EpiCache: Episodic KV Cache Management for Long Conversational Question Answering

Abstract
[…]
1 Introduction
Large language models (LLMs) (Brown et al., 2020; Yang et al., 2025; Touvron et al., 2023; Jiang et al., 2023) have significantly extended their context lengths, with LLM-based chat assistants now capable of processing hundreds of thousands to millions of tokens (Reid et al., 2024; Meta, 2025). This capability enables assistants to leverage extensive dialogue histories when generating responses, producing personalized and contextually coherent outputs (OpenAI, 2024; Anthropic, 2024), which are central requirements for real-world conversational AI applications.
Long Conversational Question Answering (LongConvQA) is the task of answering a sequence of user questions grounded in extended interaction histories, where sessions span hundreds of turns and multi-session interactions unfold over days or weeks. Recent work has formalized LongConvQA in both human-human conversations and user-AI assistant interactions (Maharana et al., 2024; Lee et al., 2025; Wu et al., 2025a). While most current systems rely on external memory modules (Zhong et al., 2024; Chhikara et al., 2025), the challenge of sustaining such long contexts under strict memory budgets remains largely unaddressed.
To bridge this gap, we study how to enable LongConvQA under constrained memory by designing a Key-Value (KV) cache control framework. The KV cache stores the Key and Value states of each token for reuse in auto-regressive generation, but its size grows linearly with context length, creating severe challenges in extended conversations. For instance, in multi-day dialogues between user and assistant (Wu et al., 2025a), the KV cache of LLaMA3.2-3B exceeds 7GB after only 30 sessions—larger than the size of the model parameters. This underscores the importance of cache management for deploying conversational AI system under resource-constrained environments.
Prior work has attempted to mitigate the growing memory footprint of KV caches in long conversations through various compression techniques (Zhang et al., 2023; Li et al., 2024; Cai et al., 2025). Yet two major limitations remain under memory-constrained settings. First, most methods apply compression after full context prefill (post-prefill), causing peak memory usage to scale linearly with input length. Second, query-dependent eviction (Li et al., 2024) optimizes for the current question but ties the cache closely to it, neglecting information needed for future queries and degrading accuracy in multi-turn conversations.
We propose EpiCache, a training-free KV cache management framework that enforces a constant memory footprint through block-wise prefill. After each block, less critical KV entries are evicted, ensuring memory consumption remains bounded. Beyond block prefill, EpiCache incorporates an episodic clustering method inspired by conversation segmentation studies (Joty et al., 2013; Galley et al., 2003). Specifically, we apply semantic clustering to group conversation history into coherent episodes, and perform episodic KV cache compression, yielding topic-specific caches within the fixed budget.
[…]
2 Background
We begin by formalizing Long Conversational Qustion-Answering (LongConvQA) in Section 2.1. We then discuss memory-constrained KV cache management in Section 2.2, comparing post- and block-prefill eviction and discussing the resulting accuracy-memory trade-offs. Finally, in Section 2.3, we review attention-guided cache compression with patched prompts and present analyses that motivate our method, EpiCache.
2.1 Long Conversational QA Formulation
We formalize LongConvQA as answering a sequence of user queries $\mathcal{Q}=\{q_{1},\dots,q_{N_{q}}\}$, with $N_{q}$ denoting the total number of queries, conditioned on long conversational history (Maharana et al., 2024; Lee et al., 2025; Wu et al., 2025a). Let the dialogue history be represented as an ordered sequence of $N_{u}$ utterances where each utterance $u_{j}$ pairs a role $r_{j}$ with text $t_{j}$:
$\mathcal{H}=\{u_{1},u_{2},\dots,u_{N_{u}}\},\quad u_{j}=(r_{j},t_{j}),\ r_{j}\in\{\text{speaker}_{1},\text{speaker}_{2}\},$ (1)
Given a long conversation $\mathcal{H}$, an LLM encodes it into a Key-Value (KV) cache $\mathrm{KV}_{\mathcal{H}}$. For $L$ layers and $H$ KV heads, encoding $N$ tokens produces $L\times H\times N$ KV entries, growing linearly with the conversation length. In this work, we focus on token-level cache compression, where KV entries of less important tokens are evicted; the resulting compressed cache is denoted as $\widetilde{\mathrm{KV}}_{\mathcal{H}}\subseteq\mathrm{KV}_{\mathcal{H}}$, and we use the terms compression and eviction interchangeably.
Our goal is to generate accurate answers for all queries $q_{1},\dots,q_{N_{q}}$} grounded in the dialogue history $\mathcal{H}$, using a compressed cache $\widetilde{\mathrm{KV}}_{\mathcal{H}}$ that satisfies a memory budget $M$, i.e., with size $L\times H\times M$, while preserving answers comparable to full KV cache ($\mathrm{KV}_{\mathcal{H}}$) based generation:
$f_{\text{LM}}(q_{i}\mid\widetilde{\mathrm{KV}}_{\mathcal{H}})\approx f_{\text{LM}}(q_{i}\mid\mathrm{KV}_{\mathcal{H}}),\quad i=1,\dots,N_{q}.$ (2)
This formulation serves as an evaluation of multi-turn conversational accuracy, where multiple query-answer pairs are grounded in the same dialogue history.
(a) Post Prefill Eviction
(b) Block Prefill Eviction
(c) Top: Peak GPU Memory Bottom: LongConvQA Score
Figure 1: KV Cache Management Analysis. (a) Post prefill eviction: eviction after full-context prefill, reducing KV size at decoding but causing unbounded memory usage. (b) Block prefill eviction: input processed in 3-token blocks with patched prompts for scoring, then evicted to 1 token. (c) Top: Peak GPU memory vs. input length on LLaMA-3.2-3B with A100. Bottom: LongConvQA accuracy of KV compression methods under post vs. block prefill on LLaMA-3.2-3B.
2.2 KV Cache Management: Post Prefill vs Block Prefill
Most existing KV compression approaches reduce cache size in the decoding stage by performing eviction after the full context has been prefilled, i.e., Post Prefill Eviction (Li et al., 2024; Feng et al., 2024; Cai et al., 2025; Kim et al., 2025). As shown in Figure 1(a), this design causes peak memory usage to grow linearly with input length, since the entire context must be cached before any eviction takes place. With optimized attention kernels (Dao, 2024), the prefill stage remains unbounded in memory demand, as observed in Figure 1(c) top.
To bound memory growth, Block Prefill Eviction (Kim et al., 2024; Corallo & Papotti, 2024; Park et al., 2025) processes the input in a block-wise way, handling one segment at a time under a fixed budget $M$. Each step adds $M_{\text{block}}$ tokens, after which eviction reduces KV cache entries back to $M$. For example, in Figure 1(b), the budget is $M=1$, and each block adds $M_{\text{block}}=3$ tokens that are evicted to $M$. This design ensures the cache never exceeds $M+M_{\text{block}}$, keeping peak GPU memory usage flat with input length as highlighted in Figure 1(c) top.
However, this bounded memory comes with a steep accuracy trade-off: when the state-of-the-art post prefill eviction method KVzip, (Kim et al., 2025; NVIDIA, 2025) is applied in the block prefill setting, LongConvQA scores (Maharana et al., 2024) degrade sharply across all budget levels. This underscores a central challenge—while block prefill guarantees constant memory usage, adapting post-prefill eviction methods to this setting severely undermines answer quality in LongConvQA.
2.3 Attention-guided KV Cache Compression
To address the accuracy degradation of block prefill eviction, prior work employs attention-based token scoring with patched prompt. Here, token importance is quantified by the cross-attention it receives from query tokens: $\mathrm{Attn}(x_{t}\!\rightarrow\!x_{i})$ which denotes the attention weight from a query token $x_{t}$ to a key token $x_{i}$. Tokens that receive higher attention from queries are considered more important, while those with lower scores are evicted to satisfy the memory budget $M$.
To provide guidance for cache eviction, the patched prompt strategy (Kim et al., 2024; Bhaskar et al., 2025) appends an auxiliary prompt of length $p$ after each block ending at token $n$. These queries $x_{n+1},\dots,x_{n+p}$} attend back to the preceding block tokens $x_{1},\dots,x_{n}$}, as illustrated in Figure 1(b). The resulting importance score $s_{i}$ of token $i$ is aggregated either by averaging or by taking the maximum, as defined in Equation 3 (the patched prompt is only used for scoring and not retained in the KV cache; all experiments in this paper adopt the maximum aggregation).
$s_{i}^{\text{avg}}=\frac{1}{p}\sum_{t=n+1}^{n+p}\mathrm{Attn}(x_{t}\!\rightarrow\!x_{i}),\quad s_{i}^{\text{max}}=\max_{t\in[n+1,n+p]}\mathrm{Attn}(x_{t}\!\rightarrow\!x_{i}).$ (3)
In block prefill, accuracy is highly dependent on the content of the patched prompt. To study this, Figure 2 presents a controlled experiment where we assume oracle access to the future user query, with inserting it as the patched prompt yielding the highest accuracy (Exact-Question) [^1] This strategy is infeasible in LongConvQA since (i) the future queries are unknown at compression time, and (ii) each new query requires a new prefill and compression step (Kim et al., 2025).. Since the dialogue history $\mathcal{H}$ in Equation 1 consists of question-answer turns, it offers an opportunity to approximate the future query with semantically related utterances. To test this idea, we embed both user queries $q_{1},\dots,q_{N_{q}}$ and conversation utterances $u_{1},\dots,u_{N_{u}}$ into a shared embedding space, compute semantic similarity scores, and construct patched prompts using the top-ranked segments.
Figure 2: Patched-prompt analysis: LoCoMo results with LLaMA3.1-8B under block prefill. Patched prompts are formed by selecting the top 10%–90% similar conversation utterances to $q_{i}$.
As expected, the results confirm that conversation utterances semantically aligned with the future question can serve as effective proxies, with the closest 10% achieving accuracy nearly matching the Exact-Question case in Figure 2. Accuracy then declines as the selected segments become less related, showing that the degree of semantic alignment directly determines the quality of answers generated from the compressed KV cache. These findings narrow our objective: during block prefill eviction, the central challenge is to identify a patched prompt that best approximates unseen future questions. To this end, we employ unsupervised methods to discover dialogue segments aligned with future queries, motivating the clustering-based approach that will be introduced in Section 3.1.
3 Method
3.1 Episodic KV Cache Management with Conversation Clustering
EpiCache consists of three stages as illustrated in Figure 3. Stage 1 (Clustering) is performed offline, where the conversation history is clustered into topical episodes, and a representative segment is identified for each episode. Stage 2 (Prefilling) leverages representative segments as a patched prompt to guide block prefill eviction, building episode-specific KV cache while keeping GPU memory bounded. Stage 3 (Decoding) embeds the incoming query, matches it to the most relevant episode, and retrieves the corresponding episodic KV cache for answer generation.
Stage 1. Conversation Clustering and Selecting Medoids.
For clustering conversation, we first divide the dialogue history $\mathcal{H}$ into segments of $w_{\text{embed}}$ utterances, denoted as $\mathcal{H}={S_{1},\dots,S_{K}}$.
$S_{k}={u_{(k-1)w_{\text{embed}}+1},\dots,u_{\min(kw_{\text{embed}},N_{u})}},\quad k=1,\dots,K,\quad K=\left\lceil\tfrac{N_{u}}{w_{\text{embed}}}\right\rceil$ (4)
Each segment $S_{k}$ is encoded with a sentence encoders (Reimers & Gurevych, 2019) $f_{\text{embed}}$ into a semantic vector representation $\mathbf{e}_{k}\in\mathbb{R}^{d}$, which captures the conversational content of the segment. We then apply K-Means clustering $\mathcal{C}(\cdot)$ to the embeddings $\{\mathbf{e}_{k}\}_{k=1}^{K}$, as illustrated in Figure 3(a):
$\mathcal{C}(\{\mathbf{e}_{k}\}_{k=1}^{K})\to\{\mathcal{E}_{1},\dots,\mathcal{E}_{E}\},\quad\bigcup_{e=1}^{E}\mathcal{E}_{e}=\{e_{1},\dots,e_{K}\}.$ (5)
This procedure partitions $\mathcal{H}$ into $E$ semantically coherent topical episodes (Raedt et al., 2024). Examples of clustering results and episode-level utterance samples are shown in Figure A3.
(a) Conversation Clustering
(b) Build Episodic KV Cache
(c) Episodic KVs based Decoding
Figure 3: EpiCache Overview. (a) offline segmentation and embedding of the conversation, followed by clustering into topical episodes. (b) Building episodic KV caches under a fixed GPU memory usage based on representative segments of each cluster. (c) an incoming query is embedded, matched to the closest episode, and the corresponding cache is retrieved for answer generation.
For each cluster $\mathcal{E}_{e}$, we can compute its centroid embedding:
$\mathbf{c}_{e}=\frac{1}{|\mathcal{E}_{e}|}\sum_{S_{k}\in\mathcal{E}_{e}}\mathbf{e}_{k},\quad S_{\text{medoid}}^{(e)}=\arg\max_{S_{k}\in\mathcal{E}_{e}}\cos(\mathbf{e}_{k},\mathbf{c}_{e}).$ (6)
We then identify the medoid segment—i.e., the conversation segment in each clusters whose embedding is most similar to the centroid in terms of semantic similarity. The medoid segment, as the representative of the cluster, contains multiple turns from both speakers and is used as the patched prompt in the subsequent block prefill eviction step.
Stage 2. Episodic KV Cache Compression.
As discussed in Section 2.3, patched prompts guide cache eviction toward retaining tokens relevant to the prompt content. Building on this insight, EpiCache uses the medoid segment of each episode as the patched prompt, enabling it to collect episode-specific KV entries under the memory budget. For each episode $e\in\{1,\dots,E\}$, we perform block prefill eviction over entire context, appending its patched prompt $\mathcal{P}_{e}$ after each block (Figure 3(b)). Attention scores are then computed with $\mathcal{P}_{e}$, and the top $M$ tokens are retained to form an episode-specific cache $C_{\mathrm{KV}}^{(e)}$. Finally, all episodic caches are collected into $\mathbb{B}=\{C_{\mathrm{KV}}^{(1)},\dots,C_{\mathrm{KV}}^{(E)}\}$ and stored offline for later retrieval.
Stage 3. Query-KVs Matching and Decoding.
At decoding time, each user query $q_{i}$ is embedded with the same encoder $f_{\text{embed}}$ used in clustering, ensuring that it lies in the same representation space as the episode centroids. The query is then matched to the closest centroid as follows:
$\mathbf{q}_{i}=f_{\text{embed}}(q_{i}),\quad e^{\dagger}=\arg\max_{e\in[1,E]}\cos(\mathbf{q}_{i},\mathbf{c}_{e}).$ (7)
As illustrated in Figure 3(c), the framework retrieves the corresponding episodic cache $C_{\mathrm{KV}}^{(e^{\dagger})}$ from $\mathbb{B}$ and conditions generation on it: $f_{\text{LM}}(q_{i}\mid C_{\mathrm{KV}}^{(e^{\dagger})},)$. This design enables query-specific retrieval from compressed episodic caches while keeping cache size bounded under memory budget $M$. This query-to-episode matching process incurs overhead from embedding, centroid matching, and cache retrieval, which will be analyzed in Section 4.4.
(a) Key similarity across layers
(b) $\Delta$ KL divergence with Full KV
(c) LongConvQA accuracy
Figure 4: Layer-wise Sensitivity Analysis and KV Budget Allocation. (a) Key state cosine similarity across normalized layer positions. (b) KL divergence is measured between block prefill (M=4M{=}4K) and full KV answer predictions, with uniform allocation as the baseline. Per-sample divergence shifts are shown when applying three allocation strategies—sensitivity-aware, PyramidKV, and retrieval head profiling-on the Realtalk benchmark with Qwen2.5-7B. (c) LongMemEval accuracy comparison across allocation methods using Qwen2.5-7B.
3.2 Sensitivity-aware Layer-wise KV Budget Allocation
We further address the accuracy degradation of block prefill by proposing KV cache budget allocation strategy. The key idea is to measure how much each layer’s Key state representation deviates under block prefill and to distribute a KV budget across layers in proportion to this deviation.
Simulating Block Prefill via Custom Masking.
To quantify the deviation caused by block prefill eviction, we introduce a custom masking scheme. Each transformer layer is represented as a simplified function $f$ that takes the previous layer’s output $X^{(\ell-1)}\in\mathbb{R}^{N\times d}$, where $N$ is the sequence length and $d$ the hidden dimension, and produces the $l_{\text{th}}$ layer output $X^{\ell}$ under a given mask $\mathcal{M}$: $X^{\ell}=f\left(X^{\ell-1},\mathcal{M}\right).$
While $\mathcal{M}$ is the causal mask, we replace it with a custom mask $\mathcal{M}^{\prime}$ that uses a budget $M$, while attending sink tokens and the most recent tokens (Xiao et al., 2024). This simulates block prefill eviction in a single forward pass, enabling direct measurement of its effect on layer representations.
Layer Sensitivity Guided KV Budget Allocation.
We quantify per-layer impact by comparing Key states produced under the causal mask $\mathcal{M}$ and the custom mask $\mathcal{M}^{\prime}$. [^2] Further details regarding the rationale for using Key states deviation are provided in Section C.1. For each layer $\ell$, the forward pass under each mask produces:
$K^{\ell}_{\text{full}}=f\left(X_{\text{full}}^{\ell-1},\mathcal{M}\right)W_{K}^{\ell},\qquad K^{\ell}_{\text{block}}=f\left(X_{\text{block}}^{\ell-1},\mathcal{M}^{\prime}\right)W_{K}^{\ell},$ (8)
where $K^{\ell}_{\text{full}}$ and $K^{\ell}_{\text{block}}$ are the $l$-th layer Key states computed under $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively. We then define layer sensitivity as the average cosine similarity between the two sets of Key vectors across heads and tokens:
$\sigma_{\ell}=\tfrac{1}{HN}\sum_{h=1}^{H}\sum_{i=1}^{N}\cos\!\big(k^{(\ell,h)}_{\text{full},i},\,k^{(\ell,h)}_{\text{block},i}\big),\quad s_{\ell}=1-\sigma_{\ell}$ (9)
Empirically, $\sigma_{\ell}$ exhibits large variation across layers yet remains consistent across different inputs in Figure 4(a) (the shadowed regions denote input variance), indicating that sensitivity is model-dependent rather than input-dependent. We define $s_{\ell}$ as the sensitivity score for layer $\ell$.
Based on this observation, we propose a sensitivity-aware budget allocation strategy that assigns larger cache budgets to layers more sensitive to block prefill and smaller budgets to less sensitive ones. Specifically, we redistribute the global budget ($M\cdot L$) according to layer sensitivity scores $s_{\ell}$, with $\alpha$ controlling how sharply the allocation emphasizes sensitive layers:
$M_{\ell}^{\text{alloc}}=\frac{s_{\ell}^{\,\alpha}}{\sum_{j=1}^{L}s_{j}^{\,\alpha}}\cdot(L\cdot M),\qquad\sum_{\ell=1}^{L}M_{\ell}^{\text{alloc}}=L\cdot M,$ (10)
We evaluate this approach by measuring how much budget allocation shifts the KL divergence between block prefill and full KV cache answer predictions, where negative values indicate closer alignment to full KV cache answer generation.
As shown in Figure 4(b), sensitivity-aware allocation shifts KL divergence by –0.80 relative to uniform allocation. In contrast, prior allocation strategies such as PyramidKV (Cai et al., 2025), which follows a pyramid-shaped budgeting, and retrieval head profiling based allocation (Wu et al., 2025b) tend to increase KL divergence under block prefill. This gap is directly reflected in task performance: as shown in Figure 4(c), sensitivity-aware allocation consistently improves LongConvQA accuracy and complements episodic cache management framework, while other allocation strategies severely degrade accuracy.
Figure 5: LongConvQA Evaluation Results (Realtalk, LoCoMo, and LongMemEval) results with fixed KV cache budget size-$M$ across four LLMs. The number of episodes (clusters) fixed to E=4E{=}4 in all experiments. The average full KV lengths of the three benchmarks are 26K, 21K, and 21K.
4 Experiments
4.1 Setup
Models and Benchmarks.
We evaluate on four pretrained LLMs: LLaMA-3.2-3B, LLaMA-3.1-8B (Grattafiori et al., 2024), Qwen2.5-3B, and Qwen2.5-7B (Qwen et al., 2025). All experiments follow the LongConvQA setup in Equation 2, where models answer various queries grounded in long conversation histories with compressed KV cache. We use three benchmarks: Realtalk (Lee et al., 2025) and LoCoMo (Maharana et al., 2024), containing multi-day human-human dialogues, and LongMemEval (Wu et al., 2025a), consisting of multi-session user-LLM conversations. Further details of LongConvQA benchmarks are provided in Section A.1.
Baselines.
We compare against comprehensive KV cache compression methods adapted to block prefill. StreamingLLM (Xiao et al., 2024) applies static retention of sink and recent tokens, SnapKV, InfiniPot, and KVzip (Li et al., 2024; Kim et al., 2024; 2025) use patched prompt based attention scoring, and KeyDiff (Park et al., 2025) scores tokens based on similarity among Key states and retains those with distinctive representations. Detailed baseline setting can be found in Section A.2.
EpiCache Setup.
EpiCache clusters the conversation history $\mathcal{H}$ into episodic segments and applies patched-prompt KV cache compression, with per-layer budgets allocated by sensitivity measurements. We use E=4E{=}4 episodes and Qwen3-0.6B (Zhang et al., 2025) as the embedding model. Layer sensitivities are profiled once on a BookSum (Kryściński et al., 2022) sample for each LLM and reused across all experiments, with the sharpness hyper-parameter set to $\alpha{=}1.1$ for the LLaMA series and $\alpha{=}1.3$ for the Qwen series. The overall process is detailed in Algorithm 1, while further setup specifications are provided in Section A.3.
4.2 Main Evaluation Results
LongConvQA Evaluation.
Figure 5 shows results on the three LongConvQA benchmarks—Realtalk, LoCoMo, and LongMemEval—under block prefill eviction with varying cache budget $M$. StreamingLLM consistently shows the lowest accuracy, while patched-prompt methods (SnapKV, InfiniPot, KVzip) and Key states similarity-based KeyDiff suffer significant accuracy degradation compared to full KV. EpiCache achieves significantly higher scores across all models and benchmakrs, exceeding baselines by around 20 scores under tightly compressed budgets (M=M{=}2-4K), and approaching full KV performance at budgets (M=M{=}6-8K). These consistent improvements across the three benchmarks highlight the effectiveness of episodic KV cache compression, along with sensitivity-aware budget allocation, in preserving conversational context under fixed memory constraints. Detailed results including sub-task accuracy are provided in Appendix D.
Figure 6: Memory Scalability up to 100K Context. Conversation histories between user and LLMs scaled to 100K tokens across four LLMs with LongMemEval. Comparison of InfiniPot and KVzip (M=6M{=}6K) with EpiCache (4 episodes, M=6M{=}6K–24K).
Memory Scalability Evaluation.
Figure 6 evaluates LongConvQA under extended conversation lengths, scaling up to 100K tokens. [^3] LongMemEval (Wu et al., 2025a) supports stacking conversation sessions with associated QA pairs, allowing conversation histories to be constructed at custom lengths. Open-source LLMs exhibit declining QA performance as context length grows to 100K, as reported in Wu et al. (2025a), and the performance gap between full KV and cache compression methods (KVzip, InfiniPot) becomes increasingly pronounced. EpiCache delivers higher accuracy than baselines at the same memory budget across all context lengths, and as the KV cache budget increases, its accuracy steadily approaches full KV, demonstrating the memory scalability of our approach.
4.3 Ablation Study
We conduct ablation studies to examine the design choices of EpiCache. The detailed setups and results are provided in Appendix B, and we summarize the design aspects here:
  • Alternative design: replacing episodic caches with a RAG-like approach that directly inputs clustered conversation segments with the query.
  • EpiCache ablation studies: effect of block size $M_{\text{block}}$; window size $w_{\text{embed}}$; encoder choice $f_{\text{embed}}$; number of episodes $E$; number of medoids for patched prompts.
4.4 Efficiency Analysis
Figure 7(a) reports decoding latency breakdown and peak GPU memory usage, comparing full KV with EpiCache under cache budgets of 2K-8K. Latency is evaluated per turn, consisting of answer decoding [^4] For fair comparison, decoding latency is measured by generating up to 10 tokens per turn. and, for EpiCache, query embedding, centroid matching, and KV cache retrieval. With fewer KV entries stored, EpiCache reduces decoding latency by up to $2.4\times$ and peak GPU memory by $3.5\times$ compared to full KV. These results demonstrate that EpiCache achieves both faster decoding and substantially lower memory usage, as block prefill bounds peak memory to the fixed cache budget throughout prefill and decoding.
Episode switching analysis in Figure 7(b) shows that retrieval overhead in EpiCache arises only when the matched episode changes between turns. We measure this by clustering each conversation, embedding subsequent turns, and tracking how often the assigned episode switches. The cumulative switch counts grow sublinearly compared to the Worst Case line (episode switched every turn), indicating that consecutive turns remain within the same episode. This conversational persistence amortizes retrieval cost, and when incorporated into the per-turn latency breakdown of Figure 7(a), the effective overhead remains under 5% while maintaining robust LongConvQA accuracy. Detailed measurements information are provided in Section D.1.
(a) Decoding latency breakdown and
peak memory usage
(b) Number of episode switch in Realtalk
Figure 7: Efficiency Analysis in Multi-Turn Conversation: (a) Per-turn decoding latency and peak GPU memory for full KV (100K) and EpiCache (E=4E{=}4) with LLaMA-3.2-3B. Query Embed and Match: query encoding and centroid matching, KVs Retrieve: loading episodic cache from CPU to GPU memory. (b) Cumulative episode switches in Realtalk with E=4E{=}4, showcasing how often episodes change across turns.
5 Related Work
KV Cache Compression.
To mitigate the growing memory cost of KV caches, prior work either quantizes states into lower precision (Hooper et al., 2024; Liu et al., 2024b) or evicts less important tokens based on attention scores (Zhang et al., 2023; Li et al., 2024; Cai et al., 2025; Kim et al., 2025). As discussed in Section 2.2, most cache eviction methods follow the post-prefill approach, which requires full context prefilling and leads to unbounded peak memory. Block prefill methods such as InfiniPot (Kim et al., 2024), FINCH (Corallo & Papotti, 2024), and KeyDiff (Park et al., 2025) bound memory but suffer sharp accuracy drops in multi-turn conversation scenario.
Retrieval-based Attention
Another line of work improves decoding efficiency by selectively retrieving only the most relevant parts of the KV cache for each query token, thereby reducing the cost of attention computation. Quest (Tang et al., 2024) retrieves KV entries at the granularity of pages, while SqueezedAttention (Hooper et al., 2025) and ClusterKV (Liu et al., 2024a) clusters Key states and loads the cluster most relevant to the query. A2ATS (He et al., 2025) apply vector quantization to construct codebooks and restore Key states to determine which parts of the KV cache to retrieve.
These methods share two key limitations. First, they operate in the post-prefill regime, assuming unbounded memory usage when building the retrieval index during prefill. Second, their retrieval units—pages, clusters, or codebooks—do not align with the episodic structure of conversations, limiting their applicability to LongConvQA under strict memory budgets.
6 Conclusion
EpiCache is a framework that combines block-wise prefill with episodic clustering and sensitivity-aware budget allocation to preserve topic-relevant context under fixed memory. Across multiple LongConvQA benchmarks, EpiCache substantially outperforms existing compression methods, showing that efficient multi-turn interaction is feasible even under strict resource constraints and marking a practical step toward memory-efficient conversational AI.
Appendix A Experimental Details
A.1 Dataset
We evaluate EpiCache on three LongConvQA benchmarks: Realtalk (Lee et al., 2025), LoCoMo (Maharana et al., 2024), and LongMemEval (Wu et al., 2025a). Three benchmarks follow the LongConvQA formulation in Section 2.1, where a long conversational history $\mathcal{H}$ is provided and the model is required to answer a sequence of queries $\mathcal{Q}={q_{1},\dots,q_{N_{q}}}$ grounded in dialogue history. This formulation is intended to assess the answer accuracy of LLMs in a multi-turn conversation.
Realtalk.
Realtalk (Lee et al., 2025) is a real-world dataset of 10 long-term conversations, where pairs of participants engaged in daily messaging for 16-21 days. Unlike LLM-simulated corpora such as LoCoMo, Realtalk captures natural dialogue including typos, abbreviations, asynchronous response gaps, and consecutive messages, while also reflecting diverse emotional expressions and shifts in persona consistency.
For evaluation, the dataset provides annotated memory probing questions across three subtasks-multi-hop, temporal reasoning, and commonsense—requiring models to recall and reason over extended histories. Following the original setup, we adopt GPT-based scoring (gpt-4o-mini-2024-07-18) to assess open-ended generation.
LoCoMo.
LoCoMo (Maharana et al., 2024) is a benchmark of long-term conversations, created through a human-machine pipeline where LLM-based agents generate dialogues grounded in distinct personas and temporal event graphs, and human annotators refine them for long-range consistency. The dataset consists of 10 conversations, each spanning up to 35 sessions with around 300 turns. The QA benchmark is divided into five subtasks: (i) Single-hop, (ii) Multi-hop, (iii) Temporal reasoning, (iv) Open-domain knowledge, and (v) Adversarial. Evaluation is conducted with open-ended genetation and report F1 score.
We exclude the adversarial subtask for the following reason. This task evaluates whether a model can detect unanswerable questions by choosing between a plausible but incorrect answer and a “no such information” response. Under KV cache compression, however, models often over-predict the latter, which spuriously inflates adversarial scores. For instance, on LLaMA-3.2-3B the score is only 12.11 with full KV, but rises to 49.78 under 4K KVzip compression. Unlike the consistent degradation seen in other subtasks (e.g., temporal or multi-hop reasoning), this inflation does not indicate real gains. Since open-source models already struggle on this task (Maharana et al., 2024), including it would be misleading. We therefore omit adversarial results from our main evaluation and defer a more careful study of unanswerability detection under compression to future work.
LongMemEval.
LongMemEval (Wu et al., 2025a) benchmarks long-term memory in user-assistant interactions with five core abilities—information extraction, multi-session reasoning, temporal reasoning, knowledge updates, and abstention—through seven question types (single-session user/assistant/preference, two-hop, multi-session synthesis, knowledge update, temporal reasoning, and abstention). A key property is its length-configurable chat histories: the benchmark provides standardized settings with extremely long contexts (e.g., up to 1.5M tokens), designed as controlled stress tests of memory and retrieval mechanisms. We follow the open-ended generation setup and report F1 scores for this dataset.
To align LongMemEval with the LongConvQA formulation in Section 2.1, we utilize the custom session stacking provided by LongMemEval [^5] https://github.com/xiaowu0162/LongMemEval to build coherent long conversations from user-LLM. Using this feature, we construct evaluation sets while preserving the original distribution of all question types. Specifically, we sample QA pairs according to the benchmark’s task-type proportions, retrieve the corresponding evidence conversation sessions, and assemble them into chronologically consistent histories. We then evaluate models at context lengths of 20K, 40K, 60K, 80K, and 100K tokens. This design allows us to test KV cache compression under scalable memory budgets.
A.2 KV Cache Compression Baseline Setup
We adapt existing KV cache compression methods to the block prefill setting for a fair comparison with our approach. The baselines include both static retention and attention-based eviction strategies, as well as similarity-based selection.
StreamingLLM.
Following Xiao et al. (2024), we retain a fixed number of sink and recent tokens throughout block prefill. Specifically, we fix the number of sink tokens to 128 for all models, while the remaining budget M128M-128 is assigned to the most recent tokens.
SnapKV.
We adapt SnapKV (Li et al., 2024) to the block prefill setting, where future queries are not accessible in LongConvQA. Following the original design, we use the window tokens—in our case the last 64 tokens of each block—as the patched prompt, and then apply the scoring function in Equation 3. Tokens with the highest attention relevance to this patched prompt are retained.
KVzip.
We adapt KVzip (Kim et al., 2025) to block prefill by treating the entire block of tokens as the patched prompt. At each block boundary, we append a repetition instruction (e.g., ”Repeat the part of the previous context exactly”) followed by the full block tokens, and then apply the patched-prompt scoring method.
InfiniPot.
We adopt the InfiniPot (Kim et al., 2024) by employing a general-purpose patched prompt designed to highlight globally important content. Specifically, we append the instruction “Summarize the previous context highlighting the most important parts.” at the end of each block and compute scores according to Equation 3. This encourages selection of semantically informative tokens across the block.
KeyDiff.
KeyDiff (Park et al., 2025) is KV cache eviction method for block prefill. For each block, it constructs an anchor key by averaging the key states of all tokens, and computes the dissimilarity score of each token as the negative cosine similarity between its key state and the anchor. These scores are then used to guide eviction. Following the original implementation, we evaluate $M_{\text{block}}\in{128,512,1024,2048}$, which includes the default setting of 128, and report results using the configuration that achieved the best performance.
To ensure fairness, all attention-based methods (SnapKV, InfiniPot, KVzip, EpiCache) use the same scoring formulation from Equation 3, and all eviction methods are combined with head-wise non-uniform token selection as suggested by Feng et al. (2024). In practice, we adopt the $\max$ aggregation for $s^{\max}_{i}$ rather than averaging across heads, since empirical results consistently showed superior performance across all baselines.
Algorithm 1 EpiCache with Layer-wise Budget Allocation Pseudo Code
[…]
2:Episodic caches $\mathbb{B}=\{C_{\mathrm{KV}}^{(1)},\dots,C_{\mathrm{KV}}^{(E)}\}$, centroids $\{\mathbf{c}_{e}\}_{e=1}^{E}$, and layer budgets $\{M^{\mathrm{alloc}}_{\ell}\}_{\ell=1}^{L}$
3:Phase A: Clustering and Prefill
4:A1. Conversation Segment & Clustering (Offline)
5:Partition $\mathcal{H}$ into $K=\lceil N_{u}/w_{\text{embed}}\rceil$ segments $\{S_{k}\}_{k=1}^{K}$ and encode $\mathbf{e}_{k}=f_{\text{embed}}(S_{k})$.
6:Run K-Means Clustering on $\{\mathbf{e}_{k}\}$ to obtain $\{\mathcal{E}_{e}\}_{e=1}^{E}$.
7:for $e=1$ to $E$ do
8:  $\mathbf{c}_{e}\!\leftarrow\!\frac{1}{|\mathcal{E}_{e}|}\sum_{S_{k}\in\mathcal{E}_{e}}\mathbf{e}_{k}$; $S^{(e)}_{\mathrm{medoid}}\!\leftarrow\!\arg\max_{S_{k}\in\mathcal{E}_{e}}\cos(\mathbf{e}_{k},\mathbf{c}_{e})$
9:  Build patched prompt $\mathcal{P}_{e}$ by concatenating utterances of $S^{(e)}_{\mathrm{medoid}}$.
10:end for
11:A2. Measure layer sensitivity & allocate KV budgets.
12:for each $x\in\mathcal{B}$ do
13:  $K^{\ell}_{\mathrm{full}}(x)\!\leftarrow\!f_{\mathrm{LM}}(x,\mathcal{M})W_{K}^{\ell}$; $K^{\ell}_{\mathrm{block}}(x)\!\leftarrow\!f_{\mathrm{LM}}(x,\mathcal{M}^{\prime})W_{K}^{\ell}$    for $\ell{=}1{:}L$
14:  $\sigma_{\ell}(x)\!\leftarrow\!\frac{1}{HN}\sum_{h=1}^{H}\sum_{i=1}^{N}\cos\!\big(k^{(\ell,h)}_{\mathrm{full},i}(x),k^{(\ell,h)}_{\mathrm{block},i}(x)\big)$
15:end for
[…]
18:A3. Build episodic KV caches.
19:for $e=1$ to $E$ do
20:  Block-wise prefill over $\mathcal{H}$, appending $\mathcal{P}_{e}$ to each block of $M_{\text{block}}$ tokens.
21:  Compute scores w.r.t. $\mathcal{P}_{e}$ with Equation 3 and retain the top $M$ tokens.
22:  $C_{\mathrm{KV}}^{(e)}\leftarrow$ compressed cache for episode $e$.
23:end for
[…]
25:Phase B: Online decoding
26:For query $q_{i}$: $\mathbf{q}_{i}\!\leftarrow\!f_{\text{embed}}(q_{i})$; $e^{\dagger}\!\leftarrow\!\arg\max_{e}\cos(\mathbf{q}_{i},\mathbf{c}_{e})$
27:Retrieve $C_{\mathrm{KV}}^{(e^{\dagger})}$ and generate with compressed cache: $f_{\mathrm{LM}}(q_{i}\mid C_{\mathrm{KV}}^{(e^{\dagger})})$.
A.3 EpiCache Setup
Overall Process.
We provide the complete procedure of EpiCache in Algorithm 1. The framework consists of two phases. In Phase A, the conversation history is clustered into topical episodes and a compressed KV cache is prepared for each episode. In Phase B, online queries are answered by retrieving the most relevant episodic cache.
In Phase A1, we segment the conversation history, embed each segment, and cluster them into $E$ episodes. This step can be performed offline, and the cost of segment encoding and K-means clustering is negligible (under a minute). Each episode is represented by its centroid and a medoid segment that serves as a patched prompt. In Phase A2, we measure per-layer sensitivity by comparing Key states under full and block-prefill masks, and allocate layer-wise budgets proportionally using the sharpness hyper-parameter $\alpha$. In Phase A3, we construct episodic KV caches by performing block-wise prefill with the patched prompt appended, and then compress the resulting caches according to the allocated budgets from Phase A2. Although prefill must be repeated for every episode, the peak memory remains flat during this process (see Figure 1(c)), making it practical for constrained-memory environments.
In Phase B, when a new query arrives, we embed the query and compute its similarity to the episode centroids. The query is then routed to the most relevant episodic cache, which is loaded for decoding. If the same cache is selected as in the previous turn, no additional retrieval is required since the cache remains resident, further reducing overhead.
Detailed Settings
For segment construction, we set the embedding window size $w_{\text{embed}}$ to 4, selected from 2,4,8{2,4,8}. We observed that performance differences among these values were minor, making $w_{\text{embed}}{=}4$ a reasonable default. To cluster segments into episodes, we apply the standard K-means algorithm with k-means++ initialization (Arthur & Vassilvitskii, 2007). This offline segmentation and clustering stage completes within one minute, incurring negligible overhead.
For sensitivity-aware budget allocation, we estimate per-layer weights using a single randomly sampled conversation from the BookSum (Kryściński et al., 2022) dataset. By performing two forward passes—one with the full causal mask and one with the block-prefill mask—we measure the layer-wise deviations and compute allocation weights. Because only one sample is used, the overhead of this calibration step is negligible.
In block prefill, the cache always maintains size $M$: as conversation segments are added in blocks of $M_{\text{block}}$, the cache can temporarily grow up to $M+M_{\text{block}}$ entries, after which eviction is applied to reduce it back to $M$. A larger $M_{\text{block}}$ enables the model to cover the entire conversation more quickly but increases the temporary peak memory footprint, while a smaller $M_{\text{block}}$ lowers peak memory at the cost of slower coverage. We set $M_{\text{block}}\in{128,512,1024,2048}$ to balance this trade-off.
For the patched prompt, we use the medoid segment of size 4, selected from 4,8,12{4,8,12}. Ablation studies show that different segment sizes yield only marginal performance differences. After constructing episodic KV caches, we offload them to offline memory (e.g., CPU) to minimize GPU memory usage. During online decoding, episodic caches are retrieved from offline memory. This design enables constrained GPU memory usage while keeping retrieval overhead manageable.
Figure A1: Alternative Design Exploration: (a) Comparison with RAG-like baseline, which feeds clustered segments directly with query-based matching. (b) Sweep of block prefill size $M_{\text{block}}$. (c) Effect of segmentation window size $w_{\text{emb}}$. (d) Encoder choice for clustering and query embedding. (e) Effect of episode number $E$ in episodic clustering. (f) Number of medoid samples in patched prompts. All experiments use LLaMA3.1-8B on the Realtalk and LoCoMo benchmarks.
Appendix B Alternative Design Exploration
B.1 Comparison with RAG-like approach.
A representative alternative design is a RAG-like approach, where conversation segments are clustered and the query is used to select the most relevant clusters, which are then directly fed into the LLM to form a KV cache under a fixed budget. The key difference from block prefill eviction is that block prefill spans the entire dialogue and selects tokens in a context-holistic manner, whereas the RAG-like method isolates only the chosen clusters when constructing the KV cache. As shown in Figure A1 (a), this leads to substantially worse performance—not only below EpiCache but also below block prefill-based compression methods such as KVzip.
This result further suggests that clustering alone is insufficient for directly constructing episodic memory inputs: conversation histories segmented only by clustering remain under-contextualized. Developing more effective clustering or episode partitioning strategies tailored to retrieval thus remains an interesting direction for future work.
B.2 EpiCache Ablation Study.
Block Size.
Figure A1 (b) varies the block size $M_{\text{block}}$ used during block prefill. Accuracy differences remain minor across settings. In practice, smaller block sizes reduce peak memory but require more iterations to span the full context, whereas larger blocks incur higher memory during prefill but process the history more efficiently. This reflects a memory-throughput trade-off, and block size can be tuned according to deployment constraints.
Clustering Design.
Figures Figure A1 (c-f) examine clustering-related hyper-parameters. (i) Figure A1 (c) sweeps the segmentation window size $w_{\text{emb}}$ for utterance grouping, showing little sensitivity and indicating robustness to segmentation granularity. (ii) Figure A1 (d) varies the encoder for clustering and query embedding. The LLM’s own embedding layer yields weaker performance, while lightweight dedicated encoders (MiniLM-L6-v2, 0.06B; Qwen3-0.6B) deliver strong gains over InfiniPot. Larger encoders (Qwen3-4B) add only marginal improvements, suggesting that small encoders suffice for EpiCache. (iii) Figure A1 (e) sweeps the number of episodes $E$, where larger $E$ improves segmentation and accuracy—especially under tight budgets—at the cost of maintaining more episodic caches offline. (iv) Figure A1 (f) varies the number of medoid samples in patched prompts, with little effect, suggesting that only a few representatives per cluster are sufficient.
Appendix C Further Analysis
(a) Key states similarity
(b) Value states similarity
(c) Layer outputs L2 distance
Figure A2: Layer-wise Sensitivity Analysis. layer-wise deviation results under the full and block masks using Qwen2.5-7B on LoCoMo conversation history. Key and Value states are measured by cosine similarity, while hidden states at layer outputs are measured by L2 distance; shaded regions indicate variance across input samples.
(a) t-SNE visualization
Episode Medoid segments examples
0. Video game A: Its a game I used to play a lot … mostly play for fun …
B: I also have finished the first game … my favorite games of all time are …
1. Movie A: I love the art and I think he is an incredible director …
B: I haven’t watched movies in a while …
2. Literature A: Have you ever read a book called …
B: I’m currently reading … a multi generational family …
3. Weather A: I love the weather today its gotten warmer
B: I just got home from work … it’s raining like crazy … parts of the city are flooded
(b) Medoid samples by cluster.
Figure A3: Episodic clustering of conversation segments. (a) t-SNE visualization of conversation clustering. (Silhouette score=0.28) (b) Medoid segments illustrate coherent topics per cluster.
C.1 Block-Prefill Sensitivity Analysis
We analyze layer-wise deviations under block prefill by comparing multiple internal states—Key, Value, and layer outputs—across Transformer layers computed with the full causal mask ($\mathcal{M}$) and the block mask ($\mathcal{M}^{\prime}$). To this end, we forward LoCoMo (Maharana et al., 2024) conversation history samples under both masks and plot the resulting internal states differences across layers, as defined in Equations 8 and 9. Specifically, each plot reports cosine similarity of Key and Value states, and L2 distance of layer outputs, respectively.
We find that Value states (Figure 2(b)) exhibit consistently low similarity across layers, offering little discriminative trend. Layer outputs, measured by L2 distance (Figure 2(c)), show a monotonic error accumulation pattern rather than meaningful variation. In contrast, Key states (Figure 2(a)) provide clear differentiation across layers. This observation motivates our use of Key state deviation as the sensitivity measure for budget allocation, as discussed in Section 3.2. Further analysis of why these trends differ across Key, Value, and layer-wise output representations is left for future work.
C.2 Conversation Clustering Analysis
In this section, we provide qualitative examples of conversation clustering to illustrate how episodic structures emerge in practice. Conversation histories are divided into segments of $w_{\text{embed}}=4$ utterances, which are then embedded using Qwen3-0.6B (Zhang et al., 2025). Segment embeddings are clustered with K-Means, and the resulting clusters are visualized in two dimensions via t-SNE, as shown in Figure A3 (a).
For each cluster, we further present representative medoid segments in Figure A3 (b). These examples demonstrate that the clustering procedure consistently groups segments into coherent topical episodes, such as games, movies, literature, or weather. The medoid samples highlight the interpretability of each episode and indicate how such episode-level partitioning can serve as the basis for episodic KV cache compression.
Appendix D Detailed Experimental Results
LLaMA3.2-3B Qwen2.5-3B
Method $M$ Decode Embed. Retr. Peak Mem. KVs Decode Embed. Retr. Peak Mem. KVs
(ms) (ms) (ms) (GB) (GB) (ms) (ms) (ms) (GB) (GB)
Full KV 100K 68.9 28.4 11.6 54.1 20.7 3.8
EpiCache 2K 28.1 6.1 2.0 8.2 0.2 36.7 5.2 1.2 7.6 0.1
4K 29.6 6.0 4.0 8.4 0.5 37.7 5.5 2.2 8.8 0.2
6K 30.1 5.5 6.0 9.3 0.7 38.9 5.4 2.9 9.5 0.2
8K 32.2 5.6 7.9 11.2 0.9 42.3 5.6 3.7 10.5 0.3
Table A1: Runtime and memory comparison under full KV and block prefill across LLaMA and Qwen. Columns: $M$ (memory budget), Decode (ms), Embed. (ms), Retrieve. (ms), Peak Memory (GB), and KVs Storage (GB). Latency is averaged per turn over 100 multi-turn queries on LongMemEval with 100K context length, reported across five runs. Measurements are conducted on an NVIDIA DGX A100 system.
D.1 Efficiency Results
Measurement Setup.
We use a server equipped with 8× NVIDIA A100-40GB (PCIe) GPUs and dual Intel Xeon Platinum 8275CL CPUs, with host-device transfers over PCIe 4.0 x16. Latency is measured on LongMemEval (Wu et al., 2025a) with 100K context length, where each evaluation spans 100 multi-turn queries. We report the average per-turn latency across five runs for both LLaMA3.2-3B and Qwen2.5-3B.
Results.
Table A1 presents the runtime and memory breakdown. Full KV incurs high decoding latency and large peak GPU memory, with LLaMA reaching 68.9ms per turn and 28.4GB memory. Notably, the KV cache storage of LLaMA (11.6GB) even exceeds the model’s own parameter size, underscoring the unsustainable cost of KV caching. By contrast, Qwen2.5-3B exhibits lower KV storage (3.8GB) due to its smaller attention head configuration (Ainslie et al., 2023) (2 heads vs. 8 heads in LLaMA).
With EpiCache, both models achieve substantial efficiency gains. On LLaMA3.2-3B, decoding latency drops from 68.9ms to 28.1ms (2.4× faster) and peak memory from 28.4GB to 8.2GB (3.5× smaller). For Qwen2.5-3B, latency improves from 54.1ms to 36.7ms (1.5× faster) and memory from 20.7GB to 7.6GB (2.7× smaller). The additional cost from query embedding and cache retrieval accounts for only about 5% of per-turn latency on average, confirming that the overhead introduced by episodic caching is minimal compared to the overall efficiency benefits.
D.2 LongConvQA Subtask Results
We provide detailed subtask results corresponding to the main LongConvQA experiments highlighted in Figure 5. For Realtalk and LoCoMo, results are reported in Table A2 for LLaMA3.2-3B and LLaMA3.1-8B and in Table A3 for Qwen2.5-3B and 7B. For LongMemEval, results across all four models are reported in Table A4.
Appendix E Future Work
EpiCache demonstrates that memory-bounded episodic KV caching can substantially improve accuracy on LongConvQA while maintaining minimal latency overhead. By clustering dialogue into episodes and dynamically retrieving episode-specific caches, it provides a practical solution that balances efficiency and effectiveness in resource-constrained environments.
There remain several promising directions for future research. First, while we adopt standard embedding-based clustering for episode construction, more advanced clustering strategies specialized for conversational structure (Raedt et al., 2024) can be developed. Such methods, orthogonal to our framework, can yield more coherent episodic boundaries and further strengthen the effectiveness of episodic KV cache management.
Second, EpiCache currently operates with a fixed number of episodes. Extending it to adaptively determine the optimal number of episodes from the conversation history could improve scalability across diverse dialogue lengths and domains. In addition, incorporating KV cache quantization (Hooper et al., 2024; Liu et al., 2024b) into episodic caches would reduce cache storage and transfer costs, alleviating the movement overhead during the retrieval stage. Exploring these extensions would further enhance the practicality of episodic KV caching for long conversational scenario.
LLaMA3.2-3B LoCoMo (Full KV length: 21.8K) Realtalk (Full KV length: 26.4K)
Method $M$ Multi-hop Temporal Open-domain Single-hop Avg Multi-hop Temporal Common Avg
Full KV 36.0 15.1 13.2 54.5 40.3 39.0 30.8 38.2 35.3
2K 17.3 3.7 17.3 17.6 14.3 20.1 10.5 25.0 16.6
SnapKV 4K 23.1 6.0 11.5 28.7 21.9 28.5 14.7 30.0 22.7
(Li et al., 2024) 6K 27.3 9.9 10.6 36.8 27.8 31.7 18.2 33.2 26.0
8K 31.2 11.9 11.7 44.3 33.1 34.1 20.3 33.7 28.0
2K 15.9 7.2 10.1 15.0 13.3 22.1 9.4 24.3 16.8
InfiniPot 4K 21.3 12.0 10.0 23.7 20.0 29.2 16.0 31.9 23.8
(Kim et al., 2024) 6K 25.6 17.0 11.9 31.3 26.1 32.8 19.7 32.4 27.0
8K 28.6 20.5 11.4 36.9 30.4 35.6 24.5 36.9 31.0
2K 10.7 3.7 11.6 11.2 9.2 10.4 5.6 19.5 9.7
KeyDiff 4K 15.7 8.3 11.5 17.9 15.1 16.8 14.2 18.2 15.9
(Park et al., 2025) 6K 20.7 11.9 13.4 23.8 20.1 20.6 14.5 24.8 18.5
8K 23.7 15.7 14.7 30.2 25.0 24.8 20.7 29.7 23.7
2K 22.0 4.5 12.0 17.3 15.2 18.7 9.5 27.9 16.0
KVzip 4K 21.9 13.3 10.0 24.5 20.8 25.0 10.1 29.4 19.1
(Kim et al., 2025) 6K 28.1 16.5 12.0 34.2 28.0 26.1 15.2 36.3 22.8
8K 31.2 21.8 11.8 41.5 33.6 30.7 18.1 34.5 25.7
2K 29.3 15.9 13.7 33.1 27.6 34.7 23.0 41.3 30.5
4K 30.4 19.7 14.2 42.3 33.6 37.4 24.7 41.0 32.4
6K 33.7 22.2 12.1 46.4 36.9 39.9 25.6 44.2 34.3
EpiCache 8K 33.9 23.9 12.7 48.2 38.3 40.0 26.6 42.9 34.6
LLaMA3.1-8B LoCoMo (Full KV length: 21.8K) Realtalk (Full KV length: 26.4K)
Method $M$ Multi-hop Temporal Open-domain Single-hop Avg Multi-hop Temporal Common Avg
Full KV 43.1 22.7 17.2 67.4 50.5 49.2 55.9 48.4 52.0
2K 23.4 6.6 14.3 25.5 20.5 20.2 15.8 34.3 20.4
SnapKV 4K 30.8 10.9 13.6 37.3 29.2 28.7 26.6 44.4 30.1
(Li et al., 2024) 6K 33.3 13.8 13.3 46.8 35.3 37.3 32.2 42.3 35.8
8K 37.7 18.1 14.2 55.5 41.8 41.5 36.8 45.2 40.0
2K 15.2 12.3 10.2 19.6 16.7 21.2 14.9 35.0 20.5
InfiniPot 4K 23.0 22.1 13.9 29.5 25.8 30.0 25.8 39.1 29.5
(Kim et al., 2024) 6K 29.8 29.6 13.9 40.1 34.5 34.4 33.3 43.4 35.3
8K 33.1 33.8 13.8 45.5 38.8 38.3 38.2 43.4 39.0
2K 18.7 4.4 17.7 21.8 17.3 13.0 8.2 28.8 13.2
KeyDiff 4K 23.3 14.1 13.8 31.5 25.3 21.9 16.5 30.9 20.9
(Park et al., 2025) 6K 27.5 18.6 14.5 39.7 31.5 29.6 25.3 33.3 28.2
8K 33.0 25.8 16.4 46.2 37.7 33.0 29.9 33.3 31.7
2K 24.4 8.4 22.8 24.7 21.2 20.6 16.6 34.6 20.9
KVzip 4K 28.8 24.3 15.3 36.8 31.4 29.5 26.4 40.0 29.7
(Kim et al., 2025) 6K 32.8 31.0 13.3 47.3 39.1 36.4 33.4 41.5 35.8
8K 37.2 35.0 13.1 54.6 44.8 38.8 41.3 41.1 40.3
2K 33.2 26.3 14.2 43.5 36.3 37.9 35.2 45.2 37.8
4K 37.7 33.8 15.9 55.8 45.4 39.6 44.9 46.7 43.0
6K 40.3 36.5 17.8 60.8 49.3 42.3 50.5 35.2 46.5
EpiCache 8K 40.9 39.6 16.9 64.8 52.2 44.2 50.9 46.5 47.5
Table A2: LongConvQA (LoCoMo and Realtalk) Evaluation: Comparison of different KV cache compression methods under block-prefill with LLaMA series models.
Qwen2.5-3B LoCoMo (Full KV length: 21.9K) Realtalk (Full KV length: 26.6K)
Method $M$ Multi-hop Temporal Open-domain Single-hop Avg Multi-hop Temporal Common Avg
Full KV 33.2 22.9 12.3 49.1 38.4 32.7 28.0 39.6 31.6
2K 14.1 8.0 11.5 10.4 10.6 12.4 5.6 23.5 11.1
SnapKV 4K 17.9 12.7 11.5 16.2 15.5 15.4 9.0 27.0 14.3
(Li et al., 2024) 6K 21.3 13.1 13.7 21.1 19.0 17.9 9.3 26.2 15.3
8K 23.2 14.3 12.1 27.4 22.9 21.6 12.6 28.8 18.7
2K 12.4 16.1 12.6 8.8 11.2 8.6 6.1 26.3 10.1
InfiniPot 4K 18.2 19.7 10.2 15.7 16.7 15.4 8.9 24.3 13.9
(Kim et al., 2024) 6K 20.0 20.0 13.9 19.9 19.6 17.5 12.0 28.2 16.7
8K 23.9 18.8 13.2 25.1 22.8 23.9 11.4 31.1 19.5
2K 10.4 14.2 13.6 7.8 10.0 3.9 9.0 23.5 8.9
KeyDiff 4K 14.0 15.8 12.0 14.1 14.3 9.2 6.0 21.6 9.6
(Park et al., 2025) 6K 19.6 16.4 9.2 20.8 18.9 11.5 10.0 22.6 12.5
8K 21.6 17.3 7.4 27.8 23.2 18.1 15.1 24.4 17.7
2K 11.8 6.1 11.9 11.4 10.4 10.6 5.3 18.2 9.4
KVzip 4K 16.2 10.3 12.6 14.7 14.0 13.3 8.4 21.7 12.4
(Kim et al., 2025) 6K 19.5 12.7 11.3 19.1 17.4 15.0 10.0 27.8 14.7
8K 21.9 14.3 14.1 24.3 21.1 18.8 11.4 28.4 17.0
2K 23.6 7.6 13.4 23.8 19.8 25.6 13.2 37.7 22.0
4K 27.1 10.7 11.4 29.7 24.1 30.0 17.1 36.9 25.4
6K 27.3 13.3 10.8 37.2 28.8 32.6 19.7 36.4 27.5
EpiCache 8K 31.3 17.0 10.0 41.7 32.7 33.9 25.0 41.1 31.1
Qwen2.5-7B LoCoMo (Full KV length: 21.9K) Realtalk (Full KV length: 26.6K)
Method $M$ Multi-hop Temporal Open-domain Single-hop Avg Multi-hop Temporal Common Avg
Full KV 36.2 19.2 16.6 59.3 44.1 38.7 52.3 43.4 45.3
2K 17.6 5.9 12.6 14.7 13.3 9.8 10.5 14.2 10.8
SnapKV 4K 23.8 8.0 13.1 24.2 20.1 15.8 16.0 24.8 17.2
(Li et al., 2024) 6K 27.8 9.1 15.2 30.2 24.4 20.6 24.4 29.2 23.5
8K 30.8 13.2 14.9 39.4 30.8 24.4 25.1 33.5 26.1
2K 15.1 14.2 12.0 12.4 13.2 10.4 15.7 15.2 13.4
InfiniPot 4K 19.2 20.0 10.8 19.9 19.2 17.1 22.7 23.7 20.5
(Kim et al., 2024) 6K 23.4 23.7 15.0 26.0 24.3 21.6 30.0 25.5 25.8
8K 27.7 14.0 14.0 32.7 29.4 26.3 29.4 31.5 29.0
2K 12.8 16.7 12.4 13.6 14.0 9.8 15.0 18.4 13.3
KeyDiff 4K 18.9 22.0 13.9 21.8 20.8 12.0 19.9 26.3 19.5
(Park et al., 2025) 6K 26.3 25.2 14.3 29.8 27.2 15.7 26.1 21.1 21.1
8K 28.3 23.7 15.9 36.4 30.8 20.3 32.7 28.0 26.9
2K 14.3 13.7 11.7 12.9 13.3 12.2 10.8 23.9 13.3
KVzip 4K 19.4 16.2 13.9 20.6 19.0 17.7 16.0 29.8 18.8
(Kim et al., 2025) 6K 24.3 19.5 12.5 27.2 24.2 22.5 24.6 32.6 24.9
8K 25.6 23.8 13.0 34.8 29.5 26.8 28.7 37.7 29.3
2K 26.4 15.4 13.2 29.3 24.9 24.1 21.3 36.6 24.7
4K 29.0 20.9 14.1 38.5 31.6 31.1 29.3 37.5 31.3
6K 32.6 24.6 15.1 46.5 37.5 33.0 39.4 40.6 36.9
EpiCache 8K 32.6 28.1 15.3 52.7 41.6 33.8 46.9 43.6 41.0
Table A3: LongConvQA (LoCoMo and Realtalk) Evaluation: Comparison of different KV cache compression methods under block-prefill with Qwen series models.
Method $M$ LLaMA3.2-3B LLaMA3.1-8B
SH TH MS TR-E TR-I KU IP Avg. SH TH MS TR-E TR-I KU IP Avg.
Full KV 21K 84.6 10.0 12.5 47.9 27.1 52.3 6.2 39.4 87.2 14.1 17.6 56.5 28.5 56.1 6.3 43.1
2K 26.9 0.8 1.8 26.6 17.4 31.4 6.1 17.7 35.6 3.8 3.3 29.8 23.3 25.4 8.8 20.2
SnapKV 4K 40.0 5.3 2.4 40.2 20.4 48.5 6.5 26.1 63.3 9.7 3.9 45.8 24.4 45.8 10.8 32.4
(Li et al., 2024) 6K 54.5 5.3 15.8 37.1 23.0 58.7 6.3 33.0 65.1 8.3 9.0 49.4 30.4 50.5 11.8 35.6
8K 67.1 7.9 12.4 37.1 27.1 56.6 7.1 35.6 74.3 13.1 13.6 53.8 25.9 52.5 10.9 38.7
2K 46.2 0.8 10.8 40.5 19.1 40.7 8.9 26.3 39.8 1.8 3.4 29.1 31.7 35.8 7.6 24.1
InfiniPot 4K 48.5 7.9 12.3 33.6 18.4 52.3 7.6 29.4 62.4 7.9 4.1 53.9 23.0 46.8 9.2 32.7
(Kim et al., 2024) 6K 60.0 2.6 12.4 33.6 25.9 51.7 6.7 31.9 81.3 11.0 13.3 54.9 28.1 54.9 5.6 40.4
8K 76.0 4.4 13.3 40.7 25.0 52.9 7.6 36.2 90.3 9.1 18.8 54.0 28.5 56.8 9.4 42.2
2K 35.3 0.5 4.1 34.3 17.7 5.7 2.6 15.1 26.0 6.8 9.5 28.8 12.3 22.8 5.5 17.1
KeyDiff 4K 54.2 2.1 2.4 34.3 15.4 34.4 6.8 24.2 60.0 14.7 12.9 41.4 20.2 31.8 6.1 29.6
(Park et al., 2025) 6K 55.8 6.5 7.4 54.3 11.4 32.2 9.0 26.9 69.2 11.2 13.6 48.4 26.1 52.1 8.5 36.7
8K 56.1 2.1 6.6 37.1 25.9 37.7 7.7 27.9 70.7 11.2 18.4 46.1 30.0 50.1 5.2 37.6
2K 30.8 0.0 1.8 30.9 15.2 30.3 7.6 18.2 33.8 7.5 8.0 36.0 19.4 27.2 9.1 21.5
KVzip 4K 44.7 2.6 6.5 37.1 15.0 37.8 7.4 24.0 56.6 9.7 6.7 35.3 24.1 42.7 12.4 29.8
(Kim et al., 2025) 6K 58.1 5.3 12.4 37.1 21.3 50.8 6.2 31.3 61.3 11.8 6.9 42.7 29.4 47.5 11.3 33.6
8K 73.5 7.9 12.5 40.7 26.2 59.1 6.9 37.5 73.6 14.4 7.7 48.6 26.9 51.4 6.2 37.1
2K 73.0 10.5 7.4 40.5 21.0 50.8 6.0 34.4 72.3 16.7 3.3 46.5 24.0 56.4 10.5 37.1
4K 79.9 12.6 16.6 41.4 27.0 53.9 9.1 39.3 87.2 14.1 17.4 56.5 28.5 54.6 3.9 42.6
6K 85.0 10.0 13.4 40.7 27.2 55.1 8.3 39.6 83.8 13.8 25.4 55.7 25.3 54.9 10.7 42.9
EpiCache 8K 85.0 10.0 12.5 40.7 26.6 56.6 6.2 39.5 88.2 13.5 17.5 56.5 28.5 55.2 6.4 43.0
Method $M$ Qwen2.5-3B Qwen2.5-7B
SH TH MS TR-E TR-I KU IP Avg. SH TH MS TR-E TR-I KU IP Avg.
Full KV 21K 80.8 14.0 15.0 50.2 23.7 59.0 9.1 40.7 88.6 39.7 35.1 32.9 35.4 47.9 12.9 46.9
2K 23.5 1.2 1.3 11.9 0.1 25.6 3.8 12.7 19.9 4.9 8.3 33.1 21.5 31.1 8.9 19.3
SnapKV 4K 48.8 6.5 0.8 11.9 1.1 27.9 4.6 18.7 42.9 10.9 22.0 33.6 28.8 38.4 10.7 29.3
(Li et al., 2024) 6K 59.3 3.9 19.6 32.4 5.9 34.5 5.7 26.4 50.4 10.9 17.0 32.9 28.4 42.6 11.3 30.7
8K 56.3 11.4 20.4 29.8 7.9 41.7 8.1 28.6 67.4 17.1 24.4 32.9 31.6 42.9 14.1 36.7
2K 31.0 1.2 11.9 22.6 5.4 17.5 4.9 14.8 29.8 1.6 2.0 33.3 23.1 32.2 11.8 20.5
InfiniPot 4K 46.2 5.3 16.5 33.3 19.4 31.5 6.8 25.2 44.7 22.6 25.3 36.4 36.7 41.6 10.3 34.0
(Kim et al., 2024) 6K 55.9 4.8 10.6 32.4 25.7 40.1 6.1 28.7 66.1 26.7 32.9 36.4 40.5 48.1 11.7 41.7
8K 70.2 6.1 15.7 35.9 27.3 47.3 7.7 34.5 78.6 34.0 33.6 32.9 41.4 50.1 13.1 45.5
2K 12.6 14.0 6.3 44.1 11.1 27.6 3.3 17.3 18.1 8.2 9.4 31.1 14.3 31.5 11.7 18.4
KeyDiff 4K 30.5 11.4 11.1 34.5 18.9 33.0 5.7 22.5 43.1 5.8 9.9 37.1 20.5 32.7 10.7 24.7
(Park et al., 2025) 6K 51.6 15.3 21.1 43.7 20.5 35.4 9.6 30.4 57.5 13.0 13.9 34.2 26.8 38.0 9.9 30.8
8K 62.6 22.6 15.4 30.7 22.0 47.5 7.8 33.9 69.6 23.6 32.9 39.1 25.6 50.2 11.9 40.2
2K 18.4 1.2 1.3 13.7 10.9 22.5 4.2 11.8 21.1 4.8 2.4 22.8 19.8 27.1 8.9 16.5
KVzip 4K 39.4 6.5 5.0 23.8 13.8 25.3 6.4 19.2 45.4 14.1 22.0 31.7 23.3 29.9 11.5 27.5
(Kim et al., 2025) 6K 51.9 1.2 20.0 44.1 17.6 34.3 5.7 27.5 57.8 23.5 22.8 33.6 31.9 34.4 11.2 33.8
8K 68.4 8.8 15.0 39.5 20.7 40.9 10.0 32.5 69.6 23.5 23.4 32.9 35.0 40.3 10.6 37.7
2K 52.6 3.5 10.0 32.4 22.9 46.7 6.7 28.7 70.4 36.9 31.1 42.4 35.9 48.5 12.4 43.7
4K 74.4 14.0 11.0 46.7 17.6 55.9 9.1 37.0 83.6 38.3 29.8 37.6 49.6 41.3 12.8 46.6
6K 77.3 16.7 15.4 46.7 22.0 55.3 10.5 39.2 86.1 38.3 33.1 32.9 39.9 48.8 12.7 46.9
EpiCache 8K 77.3 16.7 15.0 46.7 22.8 55.9 10.1 39.4 86.0 38.3 33.1 37.6 40.5 47.3 12.9 47.2
Table A4: LongConvQA (LongMemEval) Evaluation: Evaluation results with Qwen and LLaMA series models under block prefill. SH = Single Hop, TH = Two Hop, MS = Multi-Session, TR-E = Temporal Reasoning (explicit), TR-I = Temporal Reasoning (implicit), KU = Knowledge Update. IP = Implicit Preference.